# AutoGUI: Scaling GUI Grounding with Automatic Functionality Annotations from LLMs

### v1

@@ 3.4 Improving Annotation Quality via LLM-Based Verification
The functionality annotations produced by the LLM probably contain incorrect, ambiguous, and hallucinated samples (See a case in Fig. 3), which probably misleads the trained VLMs and compromises evaluation accuracy. To improve dataset quality, we prompt LLMs to verify the annotations by checking whether the targeted element $e$ fulfills the intent of the annotated functionality $f$. This process presents the LLMs with the interacted element, its UI context, the UI changes induced by this element, and the functionality generated in the previous annotation process. The LLMs are then tasked with analyzing the UI content changes before predicting whether the interacted element aligns with the given functionality. If the LLMs determine that the interacted element fulfills the functionality given its UI context, the LLMs will grant a full score (An example in Fig. H). If the interacted element is considered to mismatch the functionality, this functionality can be seen as incorrect as this mismatch indicates that it may not accurately reflect the element’s actual role within the UI context.
To mitigate the potential biases in LLMs (Panickssery et al., 2024; Zheng et al., 2023; Bai et al., 2024), two different LLMs (i.e., Llama-3-70B (AI@Meta, 2024) and Mistral-7B-Instruct-v0.2 (Jiang et al., 2023)) are employed as verifiers and prompted to output 0-3 scores. The scoring process is formulated as follows:
$score=\text{LLM}(p_{\text{verify}},e,f,s_{t},s_{t+1})$ (3)
where $p_{\text{verify}}$ denotes the verification prompt (Tab. D). Only if the two scores are both 3s do we consider the functionality label correct (More details in Sec. A.7). Although this filtering approach seems stringent, we can make up the number of annotations through scaling.
Figure 4: Element functionality annotations generated by the proposed AutoGUI pipeline for both web and mobile viewpoints.
3.5 Functionality Grounding and Referring Task Generation
After rejecting, annotating, and verifying, we obtain a high-quality UI functionality dataset containing triplets of {UI screenshot, Interacted element, Functionality}. To convert this dataset into an instruction-following dataset for training and evaluation, we generate functionality grounding and referring tasks using diverse prompt templates (see Tab. E). To mitigate the difficulty of predicting absolute values for various resolutions, the coordinates of element bounding boxes are all normalized within the range $[0,99]$ (see Fig. 4 for examples).
3.6 Explore the AutoGUI Dataset
@@
The AutoGUI pipeline finally collects 22.4k trajectories, from which we select 2k grounding samples (evenly divided between web and smartphone views) as the test set and remove the trajectories to which these samples belong. Subsequently, 702k samples are randomly selected from the remaining instances to constitute the training set. The statistics of our dataset in Tab. 2 and Sec. A.1 show that our dataset covers diverse UIs and exhibits variety in lengths and functional semantics of the annotations. Moreover, our dataset presents a unique ensemble of research challenges for developing generalist web agents in real-world settings. As shown in Tab. 1 and Fig. 2, our dataset distinguishes itself from existing literature by providing functionality-rich data as well as tasks that require VLMs to discern the contextual functionalities of elements to achieve high grounding accuracy.
4 Analysis of Data Quality
This section analyzes the reliability of the proposed annotation pipeline and data quality.
@@ 4 Analysis of Data Quality
Comparison with Human Annotation To demonstrate the superiority of the proposed automatic annotation pipeline based on open-source LLMs, $N=145$ samples (99 valid and 46 invalid) are randomly selected as a testbed for comparing the annotation correctness of a trained human annotator and the pipeline. Here, correctness is defined as $Correctness=C/(N-R)$, where $C$ and $R$ denote the numbers of correctly annotated and rejected samples, respectively. The denominator subtracts the number of rejected samples as we are more interested in the percentage of correct samples after rejecting the samples considered invalid by the annotator. The authors thoroughly check the annotation results according to the three criteria in Fig. 6: 1. Context-specificity. The functionality annotations must include context-specific descriptions to ensure one-to-one mapping between the element and its annotation. 2. Appropriate details. Avoid detailing unnecessary aspects of the UIs to keep the description focused on functionality. 3. No hallucination. The annotations must not include information not grounded in the visual context of the UIs. See more details in Sec. B.1.
@@
After experimenting with three runs, Tab. 3 shows that the proposed AutoGUI pipeline achieves high correctness comparable to the trained human annotator (r6 vs. r1). Without rejection and verification (r2), AutoGUI is inferior as it cannot recognize invalid samples. Notably, simply using the rules written by the authors can improve the correctness, which is further enhanced with the LLM-aided rejector (r4 vs. r3). Moreover, utilizing the annotating LLM itself to self-verify its annotations helps AutoGUI surpass the trained annotator (r5 vs. r1). Introducing another LLM verifier (i.e., Mistral-7B-Instruct-v0.2) brings a slight increase which results from Mistral recognizing Llama-3-70B’s incorrect descriptions of how dropdown menu options work. Overall, these results justify the efficacy of the AutoGUI annotation pipeline.
Qualitatively comparing the annotation patterns of the human and AutoGUI (Fig. N), we find that AutoGUI employs the strong LLM to generate more detailed and clear annotations which would take significantly more time for the human annotator. This result suggests that the AutoGUI pipeline can lessen the burden of collecting data for training UI-VLMs.
Impact of LLM Output Uncertainty The uncertainty of LLM outputs manifests in annotation, rejection, and verification, possibly impacting the quality of the AutoGUI dataset. To evaluate this impact, we first sample 100 valid samples to test the AutoGUI pipeline for three runs. The consistency rate is 94.5%, indicating that 94.5% of the samples possess consistent annotation outcomes (i.e. correct or incorrect) across the runs. We also test the LLM-aided rejector with 46 invalid samples and find that the rejection consistency over three runs is 79.3%. This indicates that LLM uncertainty impacts this rejection process. Nevertheless, this impact is minor due to the low prevalence of invalid samples (4% of all samples) that fail the hand-written rules.
In summary, AutoGUI exhibits annotation correctness comparable to that of human annotators and LLM output uncertainty poses a minor impact on the AutoGUI annotation process.
Figure 6: The checking criteria used for comparing AutoGUI pipeline and the human annotator.
@@
Table 3: Comparing the AutoGUI and human annotator. AutoGUI with the proposed rejection and verification achieves annotation correctness comparable to trained human annotators. One LLM means Llama-3-70B and Two LLMs include Mistral-7B-Instruct-v0.2 as well.
@@ 5 Fine-Tuning Experiments
This section validates that our dataset can enhance the GUI grounding capabilities of VLMs and that the proposed functionality grounding and referring are effective fine-tuning tasks.
@@ 5.1 Experimental Settings
Training Details We select Qwen-VL-10B (Bai et al., 2023) and SliME-8B (Zhang et al., 2024b) as the base models and fine-tune them on 25k, 125k, and 702k samples of the AutoGUI training data to investigate how the AutoGUI data enhances the UI grounding capabilities of the VLMs. The models are fine-tuned on 8 A100 GPUs for one epoch. We follow SeeClick (Cheng et al., 2024) to fine-tune Qwen-VL with LoRA (Hu et al., 2022) and follow the recipe of SliME (Zhang et al., 2024b) to fine-tune it with only the visual encoder frozen (More details in Sec. B.2).
Compared VLMs We compare with both general-purpose VLMs (i.e., LLaVA series (Liu et al., 2023b; 2024a), SliME (Zhang et al., 2024b), and Qwen-VL (Bai et al., 2023)) and UI-oriented ones (i.e., Qwen2-VL (Wang et al., 2024a), SeeClick (Cheng et al., 2024), CogAgent (Hong et al., 2023)). SeeClick finetunes Qwen-VL with around 1 million data combining various data sources, including a large proportion of human-annotated UI grounding/referring samples. CogAgent is trained with a huge amount of text recognition, visual grounding, UI understanding, and publicly available text-image datasets, such as LAION-2B (Schuhmann et al., 2022). During the evaluation, we manually craft grounding prompts suitable for these VLMs.
@@ 5.2 Experimental Results and Analysis
[…]
Type Model Size FuncPred VWB EG VWB AG MoTIF RefExp ScreenSpot
General LLaVA-1.5 (Liu et al., 2023b) 7B 3.2 12.1 13.6 7.2 4.2 5.0
LLaVA-1.5 (Liu et al., 2023b) 13B 5.8 16.7 9.7 12.3 20.3 11.2
LLaVA-1.6 (Liu et al., 2024a) 34B 4.4 19.9 17.0 7.0 29.1 10.3
SliME (Zhang et al., 2024b) 8B 3.2 6.1 4.9 7.0 8.3 13.0
Qwen-VL (Bai et al., 2023) 10B 3.0 1.7 3.9 7.8 8.0 5.2
UI-VLM Qwen2-VL (Bai et al., 2023) 7B 7.8 3.9 3.9 16.7 32.4 26.1
CogAgent (Hong et al., 2023) 18B 29.3 55.7 59.2 24.7 35.0 47.4
SeeClick (Cheng et al., 2024) 10B 19.8 39.2 27.2 11.1 58.1 53.4
Finetuned Qwen-VL-AutoGUI25k 10B 14.2 12.8 12.6 10.8 12.0 19.0
Qwen-VL-AutoGUI125k 10B 25.5 23.2 29.1 11.5 14.9 32.0
Qwen-VL-AutoGUI702k 10B 43.1 38.0 32.0 15.5 23.9 38.4
Qwen-VL-AutoGUI702k 10B 50.0 56.2 45.6 21.0 51.5 54.2
Finetuned SliME-AutoGUI25k 8B 28.0 14.0 10.6 14.3 18.4 27.2
SliME-AutoGUI125k 8B 39.9 22.0 12.0 17.8 22.1 35.0
SliME-AutoGUI702k 8B 62.6 25.4 13.6 20.6 26.7 44.0
A) AutoGUI functionality annotations effectively enhance VLMs’ UI grounding capabilities and achieve scaling effects. We endeavor to show that the element functionality data autonomously collected by AutoGUI contributes to high grounding accuracy. The results in Tab. 4 demonstrate that on all benchmarks the two base models achieve progressively rising grounding accuracy as the functionality data size scales from 25k to 702k, with SliME-8B’s accuracy increasing from merely 3.2 and 13.0 to 62.6 and 44.0 on FuncPred and ScreenSpot, respectively. This increase is visualized in Fig. J showing that increasing AutoGUI data amount leads to more precise localization performance.
After fine-tuning with AutoGUI 702k data, the two base models surpass SeeClick, the strong UI-oriented VLM on FuncPred and MOTIF. We notice that the base models lag behind SeeClick and CogAgent on ScreenSpot and RefExp, as the two benchmarks contain test samples whose UIs cannot be easily recorded (e.g., Apple devices and Desktop software) as training data, causing a domain gap. Nevertheless, SliME-8B still exhibits noticeable performance improvements on ScreenSpot and RefExp when scaling up the AutoGUI data, suggesting that the AutoGUI data helps to enhance grounding accuracy on the out-of-domain tasks.
To further unleash the potential of the AutoGUI data, the base model, Qwen-VL, is finetuned with the combination of the AutoGUI and SeeClick UI-grounding data. This model becomes the new state-of-the-art on FuncPred, ScreenSpot, and VWB EG, surpassing SeeClick and CogAgent. This result suggests that our AutoGUI data can be mixed with existing UI grounding training data to foster better UI grounding capabilities.
In summary, our functionality data can endow a general VLM with stronger UI grounding ability and exhibit clear scaling effects as the data size increases.
@@
Table 5: Comparing the AutoGUI functionality annotation type with existing types. Qwen-VL is fine-tuned with the three annotation types. The results show that our functionality data leads to superior grounding accuracy compared with the naive element-HTML data and the condensed functionality annotations.
Data Size Variant FuncPred RefExp ScreenSpot
25k w/ Elem-HTML data 5.3 4.5 5.7
w/ Condensed Func. Anno. 3.8 3.0 4.8
w/ Func. Anno. (Ours full) 21.1 10.0 16.4
125k w/ Elem-HTML data 15.5 7.8 17.0
w/ Condensed Func. Anno. 14.1 11.7 23.8
w/ Func. Anno. (Ours full) 24.6 12.7 27.0
@@
B) Our functionality annotations are effective for enhancing UI grounding capabilities. To assess the effectiveness of functionality annotations, we compare this annotation type with two existing types: 1) Naive element-HTML pairs, which are directly obtained from the UI source code (Hong et al., 2023) and associate HTML code with elements in specified areas of a screenshot. Examples are shown in Fig. 2. To create these pairs, we replace the functionality annotations with the corresponding HTML code snippets recorded during trajectory collection. 2) Brief functionality descriptions that are generated by prompting GPT-4o-mini [^2] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ to condense the AutoGUI functionality annotations. For example, a full description such as ‘This element provides access to a documentation category, allowing users to explore relevant information and guides’ is shortened to ‘Documentation category access’.
@@
After experimenting with Qwen-VL (Bai et al., 2023) at the 25k and 125k scales, the results in Tab. 5 show that fine-tuning with the complete functionality annotations is superior to the other two types. Notably, our functionality annotation type yields the largest gain on the challenging FuncPred benchmark that emphasizes contextual functionality grounding. In contrast, the Elem-HTML type performs poorly due to the noise inherent in HTML code (e.g., numerous redundant tags), which reduces fine-tuning efficiency. The condensed functionality annotations are inferior, as the consensing loses details necessary for fine-grained UI understanding. In summary, the AutoGUI functionality annotations provide a clear advantage in enhancing UI grounding capabilities.
5.3 Failure Case Analysis

### v2

@@ 3.4 Improving Annotation Quality via LLM-Based Verification
The functionality annotations produced by the LLM probably contain incorrect and hallucinated samples. To improve dataset quality, we prompt LLMs to verify the annotations, inspired by works that justify the feasibility of LLM-based verification (Zheng et al., 2023; Chen et al., 2024a; Lee et al., 2023). This process presents the LLMs with the interacted element, its UI context, the UI changes induced by the interaction, and the functionality annotation generated in the previous stage. Then, the LLMs analyze the UI content changes and predict whether the interacted element aligns with the given functionality. If the LLMs determine that the interacted element fulfills the functionality given its UI context, the LLMs will grant a full score (An example in Fig. H). If not, this functionality will be seen as incorrect as this mismatch indicates that it may not accurately reflect the element’s role within the UI context.
To mitigate the potential biases in LLMs (Panickssery et al., 2024; Bai et al., 2024), two different LLMs (i.e., Llama-3-70B and Mistral-7B-Instruct-v0.2) are employed as verifiers and prompted to output 0-3 scores (This scoring range is chosen as it empirically achieves a high verification accuracy). The scoring process is formulated as follows: $score=\text{LLM}(p_{\text{verify}},e,f,s_{t},s_{t+1})$ where $p_{\text{verify}}$ denotes the verification prompt (Tab. E). Only if the two scores are both 3s do we consider the functionality annotation correct (Details in Sec. A.8). While this approach seems stringent, we can make up the number of annotations through scaling.
@@
3.5 Task Generation
After rejecting, annotating, and verifying, we obtain a high-quality UI functionality dataset containing triplets of {UI screenshot, Interacted element, Functionality}. To convert this dataset into an instruction-following dataset for training and evaluation, we generate functionality grounding and referring tasks using diverse prompt templates. The coordinates of element bounding boxes are normalized within the range $[0,999]$ (see Fig. 3).
We finally annotate 2k grounding samples as a test set and 702k as a training set (The details of ensuring no overlap between the two sets can be seen in Sec. A.1). The statistics of our dataset in Tab. 2 and Sec. A.1 show that our dataset covers diverse UIs and exhibits variety in lengths and functional semantics of the annotations.
3.6 Analysis of Data Quality
@@
Comparison with Human Annotation $N=145$ samples (99 valid and 46 invalid) are randomly selected as a testbed for comparing the annotation correctness of a trained human annotator and our pipeline. Here, correctness is defined as $Correctness=C/(N-R)$, where $C$ and $R$ denote the numbers of correctly annotated and rejected samples, respectively. The denominator subtracts the number of rejected samples as we are more interested in the percentage of correct samples after rejecting invalid samples. The authors rigorously evaluate the annotation results based on three criteria outlined in Fig. I.
Details can be found in Sec. B.1.
@@
After experimenting with three runs, Tab. 3 shows that the AutoGUI pipeline achieves high correctness comparable to the trained human annotator (r6 vs. r1). Without rejection and verification (r2), AutoGUI is inferior as it cannot recognize invalid samples. Notably, simply using the rules written by the authors can improve the correctness, which is further enhanced with the LLM-aided rejector (r4 vs. r3). Moreover, utilizing the annotating LLM itself to self-verify its annotations helps AutoGUI surpass the trained annotator (r5 vs. r1). Introducing another LLM verifier (i.e., Mistral-7B-Instruct-v0.2) brings a slight increase which results from Mistral recognizing Llama-3-70B’s incorrect descriptions of how dropdown menu options work. Overall, these results justify the efficacy of the AutoGUI annotation pipeline.
Qualitative comparison (Fig. P) shows that our pipeline generates more detailed annotations which would take more time for the human annotator.
Impact of LLM Output Uncertainty Despite LLM output uncertainty, our pipeline achieves a high annotation consistency of 94.5%. LLM uncertainty affects rejection but has a minimal overall impact due to the low prevalence of invalid samples. More experimental details in the Appendix.
@@
Table 3: Comparing AutoGUI and human annotator. AutoGUI with the proposed rejection and verification achieves correctness comparable to the trained human annotator. One LLM means Llama-3-70B and Two LLMs include Mistral-7B-Instruct-v0.2.
@@ 4 Fine-Tuning Experiments
[…]
Type Model Size FuncPred ScreenSpot ScreenSpot-v2 MoTIF VWB EG VWB AG
General GPT-4o N/A 9.8 17.8 20.4 30.5 5.6 6.8
Llama-3.2-Vision-Instruct 11B 4.9 11.7 11.6 19.7 7.0 3.9
SliME (Zhang et al., 2024b) 8B 3.2 13.0 13.4 7.0 6.1 4.9
Qwen-VL (Bai et al., 2023) 10B 3.0 5.2 5.6 7.8 1.7 3.9
Qwen2-VL (Bai et al., 2023) 7B 38.7 66.4 66.9 71.1 55.9 62.1
UI Experts CogAgent (Hong et al., 2024) 18B 29.3 47.4 52.1 45.1 55.7 59.2
SeeClick (Cheng et al., 2024) 10B 19.8 53.4 54.0 66.5 39.2 27.2
UGround-v1-7B (Gou et al., 2025) 7B 55.8 85.9 88.0 78.4 92.7 69.9
OS-ATLAS (Wu et al., 2024) 7B 52.1 82.5 84.1 78.8 82.6 71.8
Finetuned Qwen-VL-AutoGUI702k 10B 48.7
(+45.7)
41.2
(+36.0)
40.2
(+34.6)
44.0
(+36.2)
42.1
(+40.4)
35.9
(+32.0)
SliME-AutoGUI702k 8B 62.6
(+59.4)
44.0
(+31.0)
42.5
(+29.1)
44.9
(+37.9)
25.4
(+19.3)
13.6
(+8.7)
Qwen2-VL-AutoGUI702k 7B 65.0
(+26.3)
80.0
(+13.6)
83.2
(+16.3)
72.3
(+1.2)
90.3
(+34.4)
70.9
(+8.8)
SeeClick w/ AutoGUI702k 10B 50.0
(+30.2)
54.2
(+0.8)
54.7
(+0.7)
67.0
(+0.5)
56.2
(+17.0)
45.6
(+18.4)
Figure 5: Scaling effect of the AutoGUI data. The three general-purpose VLMs are fine-tuned with three scales of AutoGUI data. Using more data consistently enhances the grounding accuracy of the three models. Note that the grounding accuracy (Y-axis) is averaged over all the element grounding benchmarks.
This section validates that our dataset effectively enhances the GUI grounding capabilities of VLMs.
@@ 4.1 Experimental Settings
Evaluation Benchmarks We base our evaluation on the UI grounding benchmarks for various scenarios: FuncPred is the test split from our collected functionality dataset. This benchmark requires a model to locate the element specified by its functionality description. ScreenSpot (Cheng et al., 2024) and ScreenSpot-v2 (Wu et al., 2024) require a model to locate elements based on short instructions on mobile, desktop, and web platforms.
VisualWebBench (VWB) (Liu et al., 2024a) is a comprehensive multi-modal benchmark assessing the understanding capabilities of VLMs in web scenarios. We select the element and action grounding tasks from this benchmark.
@@
Training Details We select SliME-8B (Zhang et al., 2024b), Qwen-VL (Bai et al., 2023), and Qwen2-VL-7B (Wang et al., 2024a) as the base models and fine-tune them on 25k, 125k, and 702k samples of the AutoGUI training data to investigate how the AutoGUI data enhances their UI grounding capabilities. We fine-tune Qwen-VL and Qwen2-VL with LoRA (Hu et al., 2022) and fine-tune SliME (Zhang et al., 2024b) with only the visual encoder frozen. We also test the benefits of our dataset for a UI expert VLM, i.e., SeeClick (Cheng et al., 2024), by adding our data to its fine-tuning data. All models are fine-tuned on 8 A100 GPUs for one epoch. (More details and hyper-parameters in Sec. B.2)
Compared VLMs We compare with both general-purpose VLMs and UI expert VLMs. During the evaluation, we manually craft grounding prompts suitable for these VLMs.
@@ 4.2 Experimental Results and Analysis
Table 5: Comparing the AutoGUI functionality annotation type with existing types. Qwen-VL is separately fine-tuned with the four annotation types. Our functionality annotation leads to superior grounding accuracy.
Data Size Variant FuncPred MOTIF ScreenSpot
25k w/ Elem-HTML 5.3 11.7 5.7
w/ Condensed Func. 3.8 19.8 4.8
w/ Func. (Ours full) 21.1 22.5 16.4
125k w/ Elem-HTML 15.5 15.8 17.0
w/ Condensed Func. 14.1 23.7 23.8
w/ Func. (Ours full) 24.6 28.7 27.0
Figure 6: An example of the 2-stage planning setting used in Sec. 5 to demonstrate the potential use of AutoGUI data. The planner (a proprietary VLM, e.g., Gemini) is bad at outputting numeric coordinates when locating elements. The grounding model, finetuned with AutoGUI functionality grounding tasks, can correctly locate the task-related target element according to the expected functionality description output by the planner.
Table 6: Applying our AutoGUI dataset to 2-stage GUI agent task planning on AITW benchmark. The results show that Qwen2-VL trained with AutoGUI functionality grounding tasks can overtake the element grounding process of the proprietary models to achieve significantly higher step accuracy. Step Acc. means the percentage of correctly planned actions while Click acc. means the percentage of correctly planned click actions.
Planner Grounding Model
General
Step acc. / Click acc.
Install
Step acc. / Click acc.
Google Apps
Step acc. / Click acc.
Single
Step acc. / Click acc.
Webshopping
Step acc. / Click acc.
Avg Step acc.
GPT-4o-mini GPT-4o-mini 14.85 / 9.58 11.17 / 5.76 12.08 / 6.85 21.09 / 11.24 10.89 / 11.22 14.01
Qwen2-VL-7B SFT w/ AutoGUI 20.43 / 20.56 25.59 / 22.49 15.25 / 12.33 25.59 / 22.49 16.15 / 20.53 18.37
(+4.36)
Gemini-2.0-flash-exp Gemini-2.0-flash-exp 26.37 / 18.16 28.49 / 26.91 30.30 / 22.88 41.94 / 28.95 20.22 / 22.65 29.50
Qwen2-VL-7B SFT w/ AutoGUI 36.34 / 36.54 50.95 / 48.95 40.99 / 40.52 50.95 / 48.95 32.83 / 43.52 39.23
(+9.73)
A) AutoGUI functionality annotations effectively enhance VLMs’ UI grounding capabilities and achieve scaling effects. We endeavor to show that the element functionality data collected by AutoGUI contributes to high grounding accuracy. The results in Tab. 4 demonstrate that the base models embrace notable performance gains on all the benchmarks. The two general-purpose VLMs (Qwen-VL and SLiME), which perform poorly, witness huge performance increases after fine-tuning with AutoGUI data. Qwen2-VL fine-tuned with our data achieves high accuracy comparable to the expert models, i.e., UGround and OS-ATLAS. Interestingly, the UI expert VLM (i.e., SeeClick) also benefits from our data, with remarkable performance gains on FuncPred and VWB.
Fig. 5 shows that the three general-purpose VLMs obtain progressively rising grounding accuracy as the AutoGUI data size scales from 25k to 702k, indicating that increasing AutoGUI data amount leads to better localization performance.
In summary, our functionality data enhances VLMs UI element grounding ability and exhibits clear scaling effects as the data size increases.
@@
B) Our functionality annotations are effective for enhancing UI grounding capabilities. To assess the effectiveness of functionality annotations, we compare this annotation type with three types: 1) Naive element-HTML pairs, which are directly obtained from the UI source code (Hong et al., 2024) and associate HTML code with elements in specified areas of a screenshot. Examples are shown in Fig. 1. To create these pairs, we replace the functionality annotations with the corresponding HTML code snippets recorded during trajectory collection. 2) Brief functionality descriptions that are generated by prompting GPT-4o-mini [^2] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ to condense the AutoGUI functionality annotations. For example, a full description such as ‘This element provides access to a documentation category, allowing users to explore relevant information and guides’ is shortened to ‘Documentation category access’.
@@
After experimenting with Qwen-VL (Bai et al., 2023) at the 25k and 125k scales, the results in Tab. 5 show that fine-tuning with the complete functionality annotations is superior to the other three types. Notably, our functionality annotation type yields the largest gain on the challenging FuncPred benchmark that emphasizes contextual functionality grounding. In contrast, the Elem-HTML type performs poorly due to the noise inherent in HTML code (e.g., numerous redundant tags), which reduces fine-tuning efficiency. The condensed functionality annotations are also inferior, as the consensing loses details necessary for fine-grained element grounding. In summary, the AutoGUI functionality annotations provide a clear advantage in enhancing UI grounding capabilities.
4.3 Grounding Failure Case Analysis